OpenStudy (anonymous):

find the n th derivative of y=sinx and hence find the 9th derivative of the function.

OpenStudy (anonymous):

\[ \cos(x) = \sin(\pi/2-x) \]

OpenStudy (anonymous):

\[ \sin(\pi/2-x) = -\sin(x-\pi/2) \]

OpenStudy (anonymous):

It helps to move out the negative, because \(\frac{d}{dx} \pi/2-x = -1\) but \(\frac{d}{dx} x-\pi/2 = 1\)